The mass in g. in 0.79 mol of water (H2O) (Hint: m=n X M)

1
Expert's answer
2021-05-09T00:00:36-0400

M(H2O)=2*1+16=18(g/mol)

m(H2O)=n*M=0.79*18=14.22(g).